# Mechanical Design

### v1

@@
[…]
Mechanical Design
[…]
Reid Graves Department of Mechanical Engineering,
Carnegie Mellon University,
5000 Forbes Avenue,
Pittsburgh, PA, 15213
[…]
Carnegie Mellon University,
5000 Forbes Avenue,
Pittsburgh, PA, 15213
email: barati@cmu.edu
Airfoil Diffusion:
synthesizing airfoils for desired performance characteristics
[…]
Résumé
The design of aerodynamic shapes, such as airfoils, has traditionally required significant computational resources and relied on predefined design parameters, which limit the potential for novel shape synthesis. In this work, we introduce a data-driven methodology for airfoil generation using a diffusion model. Trained on a dataset of preexisting airfoils, our model can generate an arbitrary number of new airfoils from random vectors, which can be conditioned on specific aerodynamic performance metrics such as lift ($C_{l}$) and drag ($C_{d}$), or geometric criteria. Our results demonstrate that the diffusion model effectively produces airfoil shapes with realistic aerodynamic properties, offering substantial improvements in efficiency, flexibility, and the potential for discovering innovative airfoil designs. This approach significantly expands the design space, facilitating the synthesis of high-performance aerodynamic shapes that transcend the limitations of traditional methods.
@@ 1 Introduction
The diffusion model, once trained, allows for the generation of new airfoil shapes from random vectors. This capability enables the synthesis of new airfoils that can be selected on the basis of desired Cl and Cd values or particular geometric criteria. Our approach offers a data-driven framework that not only enhances the efficiency of the airfoil design process, but also has the potential to discover novel shapes with superior aerodynamic performance.
@@
a) Airfoil conditioning information such as the coefficient of lift ($C_{l}$), coefficient of drag ($C_{d}$), maximum thickness, and maximum camber is passed through b) conditional embedding multi-layer perceptron (MLP) layers. The embedded conditioning information is passed to e) the denoising U-Net. During training, samples are progressively distorted to approximate Gaussian noise, as shown in c). In the inference phase, a noise vector d) along conditioning value embedding are input to the U-Net e), which outputs the conditionally generated airfoil profile in d).
@@ 3 Method
Figure 2: The conditioning input, such as $C_{d}$ or $C_{l}$ is fed through several multilayer perceptron layers to generate a context vector $z$ for conditioning the diffusion model. In the forward process, the sample y coordinates are subjected to noise for $T$ timesteps, gradually approaching gaussian noise. During the backward process, the model learns to estimate the added nois at each timestep $t$ given conditioning $z$ and $y_{t}$.
@@ 3.2.3 Process for Obtaining Coefficients
    **Preparation of Input Data**: Each airfoil profile, represented by 200 points along the $x$ and $y$ coordinates, was formatted as required by NeuralFoil.
@@
    **Prediction Using NeuralFoil**: The prepared airfoil data was fed into NeuralFoil, which predicted the corresponding $C_{l}$ and $C_{d}$ values.
@@
    **Post-Processing**: The predicted $C_{l}$ and $C_{d}$ values were recorded and used for further analysis. These coefficients were essential for filtering the synthesized airfoils and evaluating their performance characteristics.
@@ 3.3.1 Forward Process
[…]
@@ 3.3.3 Airfoil Diffusion
To predict the noise added to the samples, we used the U-net architecture [23, 28, 24]. The U-net functions similarly to an autoencoder [29], where complex data is encoded into a latent representation and then decoded back. In our application, a noised sample is passed to the U-net along with timestep information. The U-net encodes the sample into a latent space and, using skip connections [28], retains information to aid in the decoding process. The output of the U-net is the predicted noise vector that was added to the input sample.
@@ 3.4 Conditional Diffusion
While standard diffusion models are effective at generating novel airfoil samples, they typically lack mechanisms to ensure that these generated samples adhere to specific geometric or performance-based criteria. To address this limitation, we incorporate a conditional diffusion model enhanced with classifier-free guidance, a technique that allows the generation process to be steered towards desired outcomes based on conditioning parameters, such as geometric features or aerodynamic performance metrics [30, 31, 6, 7].
@@ 3.4.2 Integration into the U-Net Architecture
3.4.3 Application of Classifier-Free Guidance
Classifier-free guidance is a critical component of our conditional diffusion model. Traditional approaches to conditional generation often rely on an external classifier to steer the generative process towards desired outcomes. However, this approach can be cumbersome and may introduce additional complexity and training requirements. Instead, we use classifier-free guidance, which allows the model to simultaneously learn the diffusion process with and without conditioning inputs, effectively bypassing the need for a separate classifier.
During training, the model occasionally drops the conditioning inputs and learns to generate airfoils unconditionally. By doing so, the model develops an internal understanding of the target distribution both with and without conditioning, enabling it to use this knowledge to adjust the strength of conditioning during generation. This approach provides a more flexible and efficient means of controlling the generation process, allowing us to fine-tune the balance between adhering to the conditioning criteria and exploring novel airfoil designs.
@@ 3.4.5 Summary
The conditional diffusion model developed in this work provides a powerful and flexible framework for generating airfoil profiles that meet specific geometric and aerodynamic criteria. By integrating classifier-free guidance and conditioning on performance metrics, our approach addresses the limitations of traditional diffusion models, offering a robust solution for aerodynamic shape optimization.
@@ 3.5 Shape Similarity Measurement
To evaluate the similarity between the generated airfoils and the training dataset, we utilized the Chamfer distance [32, 33], a metric commonly used in shape analysis to quantify the dissimilarity between two sets of points. The Chamfer distance is calculated by summing the average of the minimum Euclidean distances from each point in the first set to the nearest point in the second set, and vice versa. Mathematically, the Chamfer distance between two point sets $A$ and $B$ is defined as:
@@
A lower Chamfer distance indicates a higher similarity between the airfoils, while a higher distance suggests greater diversity. The distances were calculated using PyTorch on a GPU to efficiently handle the large datasets involved.
@@ 4.1 Unconditional Generation
Figure 3: Histograms of $C_{l}$, $C_{d}$, maximum camber, and maximum thickness for the UIUC and generated samples. The generated samples exhibit narrower distributions compared to the UIUC data, with their means slightly shifted to the left of the UIUC distributions.
@@ 4.2 Conditional Generation
Figure 4: a) Box plots of the percent difference between the input conditioning values for $C_{l}$ and $C_{d}$ and the corresponding evaluated values of generated profiles. Scatter plots of the relationship between input conditioning values for $C_{d}$ and $C_{l}$ in (b) and (c), respectively. Also shown are the lines of best fit, along with the visual representation of the standard deviation from the line of best fit.
The conditional diffusion model was evaluated by generating 1000 samples conditioned on specific values of $C_{l}$ and $C_{d}$. Two separate models were trained: one conditioned on $C_{d}$ and the other on $C_{l}$. The input conditioning values were linearly spaced between the minimum and maximum values of $C_{d}$ and the minimum and twice the maximum of $C_{l}$ in the UIUC dataset (Table 1) to test the model’s potential to generate samples with greater $C_{l}$ values.
Property Maximum Minimum
$C_{d}$ 0.0814 0.0038
$C_{l}$ 1.6670 -0.1701
Table 1: Min and max values of $C_{l}$ and $C_{d}$ in the UIUC dataset.
The accuracy of the generated profiles was assessed by comparing the percent difference between the input conditioning values and the corresponding evaluated values using NeuralFoil. Figure 4 a) shows boxplots of the percent differences, highlighting the model’s performance in adhering to the conditioning parameters.
The $C_{l}$-conditioned model exhibits a wider range of percent differences with a standard deviation of 13.15, indicating a greater variability in the accuracy of the generated profiles. In contrast, the $C_{d}$-conditioned model demonstrates more consistent performance, with a lower standard deviation of 0.19. This suggests that while the $C_{l}$-conditioned model is more flexible in generating a diverse set of airfoils, it may struggle to precisely match the desired $C_{l}$ values, potentially due to the higher complexity of the lift coefficient’s dependence on airfoil geometry.
Conditioning Parameter Median Standard Deviation
$C_{d}$ 0.8176 0.1960
$C_{l}$ 0.8578 13.15
Table 2: Median and standard deviation values for the box plots in Fig. 4 a). The $C_{l}$-conditioned model exhibits a broader distribution, as indicated by its greater standard deviation.
Further analysis of the relationship between the conditioned and generated values was conducted by plotting scatter plots of the input conditioning values against the NeuralFoil-evaluated results (Fig. 4 b), a)). The mean absolute error (MAE) was calculated to quantify the accuracy of the conditioning. The $C_{d}$-conditioned model achieved an MAE of 0.0018, whereas the $C_{l}$-conditioned model had a higher MAE of 0.04. The slope of the best-fit line for the $C_{d}$-conditioned model was 0.08, compared to 0.29 for the $C_{l}$-conditioned model, indicating that the $C_{d}$-conditioned model’s outputs were less sensitive to the input conditioning values. As the $C_{l}$ conditioning values exceeded the values seen in the UIUC samples, the $C_{l}$ conditioned model generations produced a greater range of $C_{l}$ evaluation values from generations, however the model still tended to produce airfoils with greater $C_{l}$ values corresponding to increasing conditioning values. This effect is likely due to passing conditioning values to the model that were not seen in training samples.
@@ 4.2.2 Novelty of Generation
Figure 6: Boxplots of the Chamfer Distance between the UIUC dataset samples and the $C_{d}$ and $C_{l}$ conditioned models.
To quantify the novelty of the generated airfoils, we calculated the Chamfer Distance between the profiles generated by the $C_{l}$ and $C_{d}$-conditioned models and the closest matching profiles from the UIUC dataset. The boxplots in Figure 6 illustrate the distribution of Chamfer Distances for both models, indicating that the generated airfoils have a significant degree of novelty, as they differ considerably from the existing profiles in the UIUC dataset.
The Chamfer Distance values for both models are generally clustered between 0.7 and 0.95, which indicates that the generated airfoils maintain a reasonable level of similarity to the training set while still exhibiting notable differences. This range also suggests that while the model is capable of generating novel airfoils, it remains somewhat anchored to the existing design space defined by the UIUC dataset.
Notably, the Chamfer Distance distributions for both models do not extend below 0.6, nor do they exceed 1.0, aligning well with the range of y-values in the training dataset, which spans from -0.21 to 0.27. This constraint likely reflects the influence of the diffusion process and conditioning, which help the model generate realistic airfoils that avoid extreme deviations from the known design space. However, the upper end of the Chamfer Distance distribution approaching 1.0 indicates that some generated airfoils are indeed pushing the boundaries of the existing dataset, suggesting the model’s potential for creating innovative designs.
The slightly higher Chamfer Distances observed in the $C_{d}$ conditioned model, compared to the $C_{l}$ conditioned model, suggest that the $C_{d}$ model may be more effective in exploring the geometric space beyond what is represented in the training dataset. This could be due to the broader range of drag characteristics that the model is conditioned on, which might encourage more variation in the resulting airfoil shapes. In contrast, the $C_{l}$ conditioned model, while still generating novel airfoils, appears to be more conservative in its exploration of new geometries, likely due to the more constrained range of lift values it is conditioned on.
Overall, these results indicate that the conditional diffusion model is capable of generating airfoils that are distinct from those in the training dataset, with the potential to introduce novel and potentially more efficient aerodynamic shapes. However, the model’s novelty is tempered by its tendency to produce airfoils that remain within a realistic and aerodynamically viable design space.
@@ 5 Conclusion
In this paper, we presented a novel diffusion-based framework for airfoil design, leveraging the UIUC airfoil database to train a model capable of generating a wide variety of airfoil profiles. By conditioning the generation process on specific aerodynamic coefficients such as lift ($C_{l}$) and drag ($C_{d}$), our model demonstrated the ability to produce airfoils that meet desired aerodynamic and geometric criteria with high fidelity.
@@
Acknowledgment
{nomenclature}
[…]
Références
  • [1] Wang, L., Zhang, H., Wang, C., Tao, J., Lan, X., Sun, G., and Feng, J., 2024, “A Review of Intelligent Airfoil Aerodynamic Optimization Methods Based on Data-Driven Advanced Models,” https://doi.org/10.3390/math12101417Mathematics, 12(10).
  • [2] Raymer, D. P., 2006, Aircraft Design: A Conceptual Approach, 4th ed., AIAA Education Series, American Institute of Aeronautics and Astronautics.
  • [3] Wang, Y., Shimada, K., and Farimani, A., 2021, “Airfoil GAN: Encoding and Synthesizing Airfoils forAerodynamic-aware Shape Optimization,” https://doi.org/10.1093/jcde/qwad046ArXiv, abs/2101.04757.
  • [4] Morris, A., Allen, C., and Rendall, T., 2008, “Wing Design by Aerodynamic and Aeroelastic Shape Optimisation,” .
  • [5] Hermans, M., Schrauwen, B., Bienstman, P., and Dambre, J., 2014, “Automated Design of Complex Dynamic Systems,” https://doi.org/10.1371/journal.pone.0086696PLoS ONE, 9.
  • [6] Jadhav, Y., Berthel, J., Hu, C., Panat, R., Beuth, J., and Farimani, A. B., 2023, “StressD: 2D Stress estimation using denoising diffusion model,” Computer Methods in Applied Mechanics and Engineering, 416, p. 116343.
  • [7] Jadhav, Y., Berthel, J., Hu, C., Panat, R., Beuth, J., and Barati Farimani, A., 2024, “Generative Lattice Units with 3D Diffusion for Inverse Design: GLU3D,” Advanced Functional Materials, p. 2404165.
  • [8] Chen, W. and Ramamurthy, A., 2021, “Deep Generative Model for Efficient 3D Airfoil Parameterization and Generation,” https://doi.org/10.2514/6.2021-1690ArXiv, abs/2101.02744.
  • [9] Xie, H., Wang, J., and Zhang, M., 2022, “Soft and Hard Constrained Parametric Generative Schemes for Encoding and Synthesizing Airfoils,” https://doi.org/10.48550/arXiv.2205.02458ArXiv, abs/2205.02458.
  • [10] Barrett, T. R., Bressloff, N., and Keane, A., 2006, “Airfoil shape design and optimization using multifidelity analysis and embedded inverse design,” https://doi.org/10.2514/1.18766AIAA Journal, 44, pp. 2051–2060.
  • [11] Zhang, Y. and Joo, S., 2020, “Optimizing Airfoil Shape Using B-Spline Curves,” https://doi.org/10.1016/j.compfluid.2020.104468Computers & Fluids, 201, p. 104468.
  • [12] Mukesh, A. and Rathakrishnan, E., 2021, “Optimization of Airfoil Shapes Using B-Spline Curves and Genetic Algorithms,” https://doi.org/10.1016/j.ast.2020.106501Aerospace Science and Technology, 110, p. 106501.
  • [13] Li, J., Zhang, M., and Martins, J. R. R. A., 2020, “Efficient Aerodynamic Shape Optimization with Deep-Learning-Based Geometric Filtering,” https://doi.org/10.2514/1.J058281AIAA Journal, 58(10), pp. 4243–4259.
  • [14] Wei, Z., Guillard, B., Bauerheim, M., Chapin, V., and Fua, P., 2023, “Latent Representation of CFD Meshes and Application to 2D Airfoil Aerodynamics,” https://doi.org/10.2514/1.J062533AIAA Journal.
  • [15] Wei, Z., Dufour, E., Pelletier, C., Fua, P., and Bauerheim, M., 2024, “DiffAirfoil: An Efficient Novel Airfoil Sampler Based on Latent Space Diffusion Model for Aerodynamic Shape Optimization,” https://infoscience.epfl.ch/server/api/core/bitstreams/9025926a-858f-422c-82d3-134bdb8cb759/contentAIAA AVIATION Forum.
  • [16] Selig, M., 1996, “UIUC airfoil data site,” Department of Aeronautical and Astronautical Engineering University of Illinois at Urbana-Champaign, http://m-selig.ae.illinois.edu/ads/coord_database.html
  • [17] Sharpe, P. D., 2021, “AeroSandbox: A Differentiable Framework for Aircraft Design Optimization,” Master’s thesis, Massachusetts Institute of Technology.
  • [18] Koning, W. J. F., Romander, E. A., Cummings, H. V., Perez Perez, B. N., and Buning, P. G., 2023, “On Improved Understanding of Airfoil Performance Evaluation Methods at Low Reynolds Numbers,” https://doi.org/10.2514/1.C037023Journal of Aircraft.
  • [19] Gracey, W., 1958, “Summary of methods of measuring angle of attack on aircraft,” National Advisory Committee for Aeronautics, https://ntrs.nasa.gov/citations/19930085167Technical Note 4351.
  • [20] Sharpe, P., 2023, “NeuralFoil: An airfoil aerodynamics analysis tool using physics-informed machine learning,” https://github.com/peterdsharpe/NeuralFoil.
  • [21] Croitoru, F.-A., Hondru, V., Ionescu, R. T., and Shah, M., 2022, “Diffusion Models in Vision: A Survey,” https://doi.org/10.1109/TPAMI.2023.3261988IEEE Transactions on Pattern Analysis and Machine Intelligence, 45, pp. 10850–10869.
  • [22] Ho, J., Saharia, C., Chan, W., Fleet, D. J., Norouzi, M., and Salimans, T., 2021, “Cascaded Diffusion Models for High Fidelity Image Generation,” J. Mach. Learn. Res., 23, pp. 47:1–47:33.
  • [23] Ho, J., Jain, A., and Abbeel, P., 2020, “Denoising Diffusion Probabilistic Models,” arXiv preprint arxiv:2006.11239.
  • [24] Ogoke, F., Liu, Q., Ajenifujah, O., Myers, A., Quirarte, G., Beuth, J., Malen, J., and Farimani, A. B., 2023, “Inexpensive High Fidelity Melt Pool Models in Additive Manufacturing Using Generative Deep Diffusion,” 2311.16168
  • [25] Yang, L., Zhang, Z., Hong, S., Xu, R., Zhao, Y., Shao, Y., Zhang, W., Yang, M.-H., and Cui, B., 2022, “Diffusion Models: A Comprehensive Survey of Methods and Applications,” https://doi.org/10.1145/3626235ACM Computing Surveys, 56, pp. 1 – 39.
  • [26] Sohl-Dickstein, J., Weiss, E. A., Maheswaranathan, N., and Ganguli, S., 2015, “Deep Unsupervised Learning using Nonequilibrium Thermodynamics,” 1503.03585
  • [27] DeGroot, M. H. and Schervish, M. J., 2014, Probability and Statistics, 4th ed., Pearson, Boston.
  • [28] Ronneberger, O., Fischer, P., and Brox, T., 2015, “U-Net: Convolutional Networks for Biomedical Image Segmentation,” 1505.04597, https://arxiv.org/abs/1505.04597
  • [29] Bank, D., Koenigstein, N., and Giryes, R., 2020, “Autoencoders,” ArXiv, abs/2003.05991.
  • [30] Ho, J. and Salimans, T., 2022, “Classifier-Free Diffusion Guidance,” 2207.12598, https://arxiv.org/abs/2207.12598
  • [31] Hu, B., Layton, A., and Chen, H., 2024, “Drug Discovery SMILES-to-Pharmacokinetics Diffusion Models with Deep Molecular Understanding,” 2408.07636, https://arxiv.org/abs/2408.07636
  • [32] Barrow, H., Tenenbaum, J., Bolles, R., and Wolf, H., 1977, “Parametric correspondence and chamfer matching: Two new techniques for image matching,” Proceedings of the International Joint Conference on Artificial Intelligence, pp. 659–663.
  • [33] Lin, C. and Zhang, H., 2023, “Hyperbolic Chamfer Distance for Point Cloud Completion,” Proceedings of the IEEE/CVF International Conference on Computer Vision (ICCV), pp. 2292–2300, https://openaccess.thecvf.com/content/ICCV2023/papers/Lin_Hyperbolic_Chamfer_Distance_for_Point_Cloud_Completion_ICCV_2023_paper.pdf
  • [34] Akram, M. T. and Kim, M.-H., 2021, “Aerodynamic Shape Optimization of NREL S809 Airfoil for Wind Turbine Blades Using Reynolds-Averaged Navier Stokes Model—Part II,” Applied Sciences, 11(5), p. 2211.

### v2

# Airfoil Diffusion: Denoising Diffusion Model For Conditional Airfoil Generation

@@
Abstract
The design of aerodynamic shapes, such as airfoils, has traditionally required significant computational resources and relied on predefined design parameters, which limit the potential for novel shape synthesis. In this work, we introduce a data-driven methodology for airfoil generation using a diffusion model. Trained on a dataset of preexisting airfoils, our model can generate an arbitrary number of new airfoils from random vectors, which can be conditioned on specific aerodynamic performance metrics such as lift ($C_{l}$) or drag ($C_{d}$). Our results demonstrate that the diffusion model effectively produces airfoil shapes with realistic aerodynamic properties, offering substantial improvements in efficiency, flexibility, and the potential for discovering innovative airfoil designs. This approach significantly expands the design space, facilitating the synthesis of high-performance aerodynamic shapes that transcend the limitations of traditional methods.
@@ Abstract
Article Highlights
  • A novel denoising diffusion model generates airfoil shapes that could improve aerodynamic performance.
  • The model can be steered towards generation of profiles with specific lift and drag characteristics.
  • This approach expands potential design techniques in aerospace applications.
@@ 1 Introduction
The diffusion model, once trained, allows for the generation of new airfoil shapes from random vectors. This capability enables the synthesis of new airfoils that can be selected on the basis of desired Cl or Cd values. Our approach offers a data-driven framework that not only enhances the efficiency of the airfoil design process, but also has the potential to discover novel shapes with superior aerodynamic performance.
@@
a) Airfoil conditioning information such as the coefficient of lift ($C_{l}$), coefficient of drag ($C_{d}$), maximum thickness, and maximum camber is passed through b) conditional embedding multi-layer perceptron (MLP) layers. The embedded conditioning information is passed to e) the denoising U-Net. During training, samples are progressively distorted to approximate Gaussian noise, as shown in c). In the inference phase, a noise vector d) along conditioning value embedding are input to the U-Net e), which outputs the conditionally generated airfoil profile in f).
@@ 3 Method
Figure 2: The conditioning input, such as $C_{d}$ or $C_{l}$ is fed through several multilayer perceptron layers to generate a context vector $z$ for conditioning the diffusion model. In the forward process, the sample y coordinates are subjected to noise for $T$ timesteps, gradually approaching gaussian noise. During the backward process, the model learns to estimate the added noise at each timestep $t$ given conditioning $z$ and $y_{t}$.
@@ 3.2.3 Process for Obtaining Coefficients
    Preparation of Input Data: Each airfoil profile, represented by 200 points along the $x$ and $y$ coordinates, was formatted as required by NeuralFoil.
@@
    Prediction Using NeuralFoil: The prepared airfoil data was fed into NeuralFoil, which predicted the corresponding $C_{l}$ and $C_{d}$ values.
@@
    Post-Processing: The predicted $C_{l}$ and $C_{d}$ values were recorded and used for further analysis. These coefficients were essential for filtering the synthesized airfoils and evaluating their performance characteristics.
@@ 3.3.1 Forward Process
$\beta_{1},\beta_{2},\dots,\beta_{T}$ $\in\left[0,1\right)$
@@ 3.3.3 Airfoil Diffusion
To predict the noise added to the samples, we used the U-net architecture [23, 28, 24]. The U-net functions similarly to an autoencoder, where complex data is encoded into a latent representation and then decoded back. In our application, a noised sample is passed to the U-net along with timestep information. The U-net encodes the sample into a latent space and, using skip connections [28], retains information to aid in the decoding process. The output of the U-net is the predicted noise vector that was added to the input sample.
@@ 3.4 Conditional Diffusion
While standard diffusion models are effective at generating novel airfoil samples, they typically lack mechanisms to ensure that these generated samples adhere to specific geometric or performance-based criteria. To address this limitation, we incorporate a conditional diffusion model that allows the generation process to be steered towards desired outcomes based on conditioning parameters, such as geometric features or aerodynamic performance metrics [29, 30, 6, 7].
@@ 3.4.4 Summary
The conditional diffusion model developed in this work provides a powerful and flexible framework for generating airfoil profiles that meet specific aerodynamic criteria. By integrating classifier-free guidance and conditioning on performance metrics, our approach addresses the limitations of traditional diffusion models, offering a robust solution for aerodynamic shape optimization.
@@ 3.5 Shape Similarity Measurement
To evaluate the similarity between the generated airfoils and the training dataset, we utilized the Chamfer distance [31, 32, 33], a metric commonly used in shape analysis to quantify the dissimilarity between two sets of points. The Chamfer distance is calculated by summing the average of the minimum Euclidean distances from each point in the first set to the nearest point in the second set, and vice versa. Mathematically, the Chamfer distance between two point sets $A$ and $B$ is defined as:
@@
A lower Chamfer distance indicates a higher similarity between the airfoils, while a higher distance suggests greater diversity. The distances were calculated using PyTorch on a GPU to efficiently handle the large quantity of computations involved.
@@ 4.1 Unconditional Generation
Figure 3: Histograms of $C_{l}$, $C_{d}$, maximum camber, and maximum thickness for the UIUC (orange) and generated (blue) samples. The generated samples exhibit narrower distributions compared to the UIUC data, with their means slightly shifted to the left of the UIUC distributions.
@@ 4.2 Conditional Generation
Figure 4: Scatter plots of the relationship between input conditioning values for $C_{d}$ and $C_{l}$ in (a) and (b), respectively. Also shown are the lines of best fit, along with the visual representation of the standard deviation from the line of best fit. In (b), the purple region depicts conditioning on values of $C_{l}$ seen in the training dataset, and the red region shows conditioning on values not seen in the UIUC dataset.
The conditional diffusion model was evaluated by generating 1000 samples conditioned on specific values of $C_{l}$ and $C_{d}$. Two separate models were trained: one conditioned on $C_{d}$ and the other on $C_{l}$. The input conditioning values were linearly spaced between the minimum and maximum values of $C_{d}$ and the minimum and twice the maximum of $C_{l}$ in the UIUC dataset to test the model’s potential to generate samples with greater $C_{l}$ values.
Analysis of the relationship between the conditioned and generated values was conducted by plotting scatter plots of the input conditioning values against the NeuralFoil-evaluated results (Fig. 4 (b), (a)). The mean absolute error (MAE) was calculated to quantify the accuracy of the conditioning. The $C_{d}$-conditioned model achieved an MAE of 0.0018, whereas the $C_{l}$-conditioned model had a higher MAE of 0.04. The slope of the best-fit line for the $C_{d}$-conditioned model was 0.07, compared to 0.25 for the $C_{l}$-conditioned model, indicating that the $C_{d}$-conditioned model’s outputs were less sensitive to the input conditioning values. As the $C_{l}$ conditioning values exceeded the values seen in the UIUC samples, the $C_{l}$ conditioned model generations produced a greater range of $C_{l}$ evaluation values from generations, however the model still tended to produce airfoils with greater $C_{l}$ values corresponding to increasing conditioning values. This effect is likely due to passing conditioning values to the model that were not seen in training samples.
@@ 4.2.2 Novelty of Generation
Figure 6: Boxplots of the Chamfer Distance between the UIUC dataset samples and the $C_{d}$ and $C_{l}$ conditioned models. Both models’ chamfer distance is predominantly around 0.005 to 0.025, with outliers as great as 0.08. The $C_{d}$ model exhibits slightly greater chamfer distances than the $C_{l}$ model.
To assess the novelty of the generated airfoils, we computed the Chamfer Distance between the profiles generated by the $C_{l}$- and $C_{d}$-conditioned models and their closest matching profiles from the UIUC airfoil dataset. As illustrated in Figure 6, the distribution of Chamfer Distances for both models highlights the extent of deviation from existing profiles in the dataset. This metric is crucial in capturing geometric dissimilarities between two sets of points, where a higher Chamfer Distance signifies a greater deviation or novelty, and a distance of zero would indicate an identical match.
To interpret the Chamfer Distance in the context of this study, we first examine the range of y-coordinates across both datasets. In the UIUC dataset, the minimum and maximum y-values range from -0.206 to 0.273, whereas for the $C_{d}$ and $C_{l}$-conditioned models, the corresponding ranges are -0.019 to 0.123 and -0.083 to 0.135, respectively. Given these ranges, we establish that a Chamfer Distance greater than 0.5 would indicate substantial novelty, as it exceeds the maximum point-to-point variation within any dataset.
For both models, Chamfer Distance values predominantly fall between 0.005 and 0.025, with outliers reaching up to 0.08. These values represent 1 to 16 percent of the maximal variation observed in the UIUC dataset, implying that even the most conservative samples exhibit at least 1 percent deviation from the training data. In contrast, the most novel generated airfoils demonstrate a geometric deviation of up to 16 percent, suggesting a meaningful degree of novelty while remaining within a physically plausible range.
The slightly elevated Chamfer Distances observed in the $C_{d}$-conditioned model relative to the $C_{l}$-conditioned model suggest that the $C_{d}$ model explores a broader geometric space. This broader exploration may stem from the wider range of drag values available in the training set, enabling the model to generate more varied and potentially unconventional shapes. Conversely, the $C_{l}$-conditioned model, while still producing novel airfoil profiles, appears more conservative, likely due to the narrower range of lift characteristics it is conditioned on.
Overall, these results confirm that the conditional diffusion models generate airfoil designs that deviate from the training dataset, introducing novel shapes. However, the generated airfoils remain within a viable aerodynamic design space, offering a balance between innovation and practicality. The higher novelty observed in the $C_{d}$ model suggests that conditioning on drag may facilitate greater exploration of the design space, making it a useful tool for discovering unique and potentially more efficient airfoil geometries.
@@ 5 Conclusion
In this paper, we presented a novel diffusion-based framework for airfoil design, leveraging the UIUC airfoil database to train a model capable of generating a wide variety of airfoil profiles. By conditioning the generation process on specific aerodynamic coefficients such as lift ($C_{l}$) and drag ($C_{d}$), our model demonstrated the ability to produce airfoils with desirable aerodynamic and geometric criteria.
@@
Declarations
  • Funding: No funding was received to assist with the preparation of this manuscript.
  • Competing interests: The authors did not receive support from any organization for the submitted work.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: All training data was collected from the UIUC airfoil database at [16].
  • Materials availability: Not applicable.
  • Code availability: Code used in this publication is available at https://github.com/gravesreid/Airfoil_Diffusion_Generator.git.
  • Author contribution:
    • Reid Graves: Methodology, Software, Data Curation, Analysis, Writing– Original Draft, Visualization, Writing– Review and Editing.
    • Amir Barati Farimani: Conceptualization, Supervision, Writing– Review and Editing.